OpenStudy (anonymous):

Which of the following points are on the line given by the equation y = 5x? Check all that apply. A. (0, 1) B. (-1, 5) C. (-1, -5) D. (3, 6) E. (4, 2) F. (3, 15)

OpenStudy (anonymous):

C and F: 0*5=0, not 1 -1*5=-5, not 5 -1*5=-5->On the line 3*5=15, not 6 4*5=20, not 2 3*5=15->On the line